40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 16977     Entry time: Thu Jul 7 18:18:19 2022     In reply to: 16929     Reply to this: 16978
Author: yuta 
Type: Update 
Category: LSC 
Subject: Actuator calibration of ETMX and ETMX 

(This is a complete restore of elog 40m/16970 from July 5, 2022 at 14:34)

ETMX and ETMY actuators were calibrated using single arm lock by taking the actuation efficiency ratio between ITMs. Below is the result.

ETMX :  2.65e-9 /f^2 m/counts (0.5007 times ITMX)
ETMY : 10.91e-9 /f^2 m/counts (2.3017 times ITMY)

Motivation:
- ETMX and ETMY actuators seemed to be unbalanced when locking DARM (see 40m/16968)

What we did:
- Reverted to C1:LSC-ETMX_GAIN = 1
- XARM was locked using POX11_I_ERR (42dB whitening gain, 132.95 deg for demod phase) with ETMX and C1:LSC-XARM_GAIN=0.06
- YARM was locked using POY11_I_ERR (18dB whitening gain, -66.00 deg for demod phase) with ETMX and C1:LSC-YARM_GAIN=0.02
- OLTFs for each was measured to be Attachment #1; UGF was ~180 Hz for XARM, ~200 Hz for YARM.
- Measured TF from C1:LSC-(E|I)TM(X|Y)_EXC to C1:LSC-(X|Y)ARM_IN1 (see Attachment #2)
- Took the ratio between ITM actuation and ETM actuation to calculate ETM actuation. For ITM actuation, we used the value measured using MICH (see 40m/16929). The average of the ratio in the frequency range 70-150 Hz was used.

Files:
- Measurement files live in https://git.ligo.org/40m/measurements/-/tree/main/LSC/XARM and YARM
- Script for calculation lives at https://git.ligo.org/40m/scripts/-/blob/main/CAL/ARM/ETMActuatorCalibration.ipynb

Discussion:
- ETMX actuation is 4.12 times less compared with ETMY. This is more or less consistent with what we measured in 40m/16968, but we didn't do loop-correction at that time.
- We should check if this imbalance is as expected or not.

Summary of actuation calibration so far:
BS   : 26.08e-9 /f^2 m/counts (see 40m/16929)
ITMX :  5.29e-9 /f^2 m/counts (see 40m/16929)
ITMY :  4.74e-9 /f^2 m/counts (see 40m/16929)
ETMX :  2.65e-9 /f^2 m/counts (0.5007 times ITMX)
ETMY : 10.91e-9 /f^2 m/counts (2.3017 times ITMY)

 

Attachment 1: Screenshot_2022-07-05_14-52-01_OLTF.png  67 kB  Uploaded Thu Jul 7 19:19:35 2022  | Show | Show all
Attachment 2: Screenshot_2022-07-05_14-54-03_TF.png  97 kB  Uploaded Thu Jul 7 19:19:43 2022  | Show | Show all
Attachment 3: Screenshot_2022-07-05_14-56-41_Ratio.png  70 kB  Uploaded Thu Jul 7 19:19:54 2022  | Show | Show all
ELOG V3.1.3-